Attempts: Average: /6 8. The effects of property rights on achieving efficiency Consider a river found in the city of Philadelphia, and then answer the questions that follow The city has a resort whose visitors use the river for recreation. The city also has a research lab that dumps industrial waste into the river. This pollutes the river and makes it a less desirable vacation destination. That is, the research lab's waste decreases the resort's economic profit. Suppose that the research lab could use a different production method that involves recycling water. This would reduce the pollution in the river to evels safe for recreation, and the resort would no longer be affected. If the research lab uses the recycling method, then the research lab's economic profit is $2,200 per week, and the resort's economic profit is $3,200 per week. If the research lab does not use the recycling method, then the research lab's economic profit is $3,000 per week, and the resort's economic profit is $2,000 per week. These figures are summarized in the following table.
Complete the following table by computing the total profit (the research lab's economic profit and the resort's economic profit combined) with and without recycling Profit Research Lab Resort Total Action (Dollars) (Dollars) (Dollars) No Recycling 3,000 2,000 Recycling 2,200 3,200 Total economic profit is highest when the recycling production method is When the research lab uses the recycling method, the resort earns $3,200-$2,000 = $1,200 more per week than it does with no recycling Therefore, the resort should be willing to pay up to $1,200 per week for the research lab to recycle water. However, the recycling method decreases the research lab's economic profit by $3,000-$2,200 = $800 per week. Therefore, the research lab should be willing to use the recycling method if it is compensated with at least $800 per week. Suppose the resort has the property rights to the river. That is, the resort has the right to a clean (unpolluted) river. In this case, assuming the two firms can bargain at no cost, the research lab will the recycling method and will pay the resort per week Now, suppose the research lab has the property rights to the river, including the right to pollute it. In this case, assuming the two firms can bargain at no cost, the research lab will the recycling method, and the resort will pay the research lab per week The research lab will make the most economic profit when True or False: The research lab will use the recycling method only if the resort has the property rights True False
Answer #1

No recycling = 3000+2000= 5000

Recycling = 2200+3200 = 5400

Total economic profit is highest when the recycling production method is used.

Blanks:-

1-Use

2-$

3-Use

4-Between $800 and $1200

The research lab will make the most economic profit when it has property rights to a clean river.

The statement is False